# JEL: C81, O32

### v1

@@ Appendix D Prompt Repository
This appendix serves as a repository for figures displaying each of the prompts considered in Section B.3. Throughout, the dummy text “{abstract}” is placed in the location where the abstract of a publication would be input. The baseline “True/False” Prompt 1.0 is displayed in Figure D.8. Refinements to this prompt are displayed in Figures D.9, D.10, and D.12. The prompt that asks the model to categorize excluded publications, i.e., Prompt 2.0, is displayed in Figure D.14. A refinement to this prompt is given in Figure D.15. The prompt that asks the model to explain why it has excluded a publication, i.e., Prompt 3.0, is displayed in Figure D.17. In turn, a refinement to this prompt is given in Figure D.18.
Figure D.8. Prompt 1.0: True/False, Base
Figure D.9. Prompt 1.1: True/False, Initial Examples
Figure D.10. Prompt 1.2: True/False, Extended Examples, Short
@@
Figure D.11. Prompt 1.2: True/False, Extended Examples, Short (Cont.)
Figure D.12. Prompt 1.3: True/False, Extended Examples, Long
@@
Figure D.13. Prompt 1.3: True/False, Extended Examples, Long (Cont.)
Figure D.14. Prompt 2.0: Categorize Exclusion Restriction
Figure D.15. Prompt 2.1: Categorize Exclusion Restriction, Examples
Figure D.16. Prompt 2.1: Categorize Exclusion Restriction, Examples (Cont.)
Figure D.17. Prompt 3.0: Provide Reason for Exclusion
Figure D.18. Prompt 3.1: Provide Reason for Exclusion, Examples
Figure D.19. Prompt 3.1: Provide Reason for Exclusion, Examples (Cont.)

### v2

@@ Appendix E Prompt Repository
This appendix serves as a repository for figures displaying each of the prompts considered in Section C.3. Throughout, the dummy text “{abstract}” is placed in the location where the abstract of a publication would be input. The baseline “True/False” Prompt 1.0 is displayed in Figure E.1. Refinements to this prompt are displayed in Figures E.2, E.3, and E.5. The prompt that asks the model to categorize excluded publications, i.e., Prompt 2.0, is displayed in Figure E.7. A refinement to this prompt is given in Figure E.8. The prompt that asks the model to explain why it has excluded a publication, i.e., Prompt 3.0, is displayed in Figure E.10. In turn, a refinement to this prompt is given in Figure E.11.
Figure E.1. Prompt 1.0: True/False, Base
Figure E.2. Prompt 1.1: True/False, Initial Examples
Figure E.3. Prompt 1.2: True/False, Extended Examples, Short
@@
Figure E.4. Prompt 1.2: True/False, Extended Examples, Short (Cont.)
Figure E.5. Prompt 1.3: True/False, Extended Examples, Long
@@
Figure E.6. Prompt 1.3: True/False, Extended Examples, Long (Cont.)
Figure E.7. Prompt 2.0: Categorize Exclusion Restriction
Figure E.8. Prompt 2.1: Categorize Exclusion Restriction, Examples
Figure E.9. Prompt 2.1: Categorize Exclusion Restriction, Examples (Cont.)
Figure E.10. Prompt 3.0: Provide Reason for Exclusion
Figure E.11. Prompt 3.1: Provide Reason for Exclusion, Examples
Figure E.12. Prompt 3.1: Provide Reason for Exclusion, Examples (Cont.)